OpenStudy (anonymous):

I need some help with IUPAC naming for the following questions (not decided, just if I need help) 2-3 questions should give me the idea of it. Any help is highly appreciated :)

OpenStudy (abb0t):

What's the question(s)?

OpenStudy (anonymous):

[Co(NH3)5 Cl]Cl2

OpenStudy (anonymous):

This question please :3 !!

OpenStudy (abb0t):

So, you have, just to be clear, \(\sf \color{red}{[Co(NH_3)_5Cl]~Cl_2}\) ?

OpenStudy (abb0t):

The rules state that the central metal atom takes on the end name. Additionally, you name them in alphabetical order, along with the appropriate number prefix. For the given example, you have: 5 ammonia, and 1 chorine, so you have pentaamminechlorocobalt (III) chloride (I think you can also use Pentaamminemonochlorocobalt(III) chloride) mono meaning 1. The roman numeral III indicates the oxidation state of the central metal atom. and the chloride is simply the Cl\(_2\) on the OUTSIDE of the brackets.

OpenStudy (anonymous):

@abb0t I have a doubt though, I am obviously going wrong with one of my facts but anyway, here's why I am not getting the same 'OXIDATION NUMBER'

OpenStudy (anonymous):

Cl has -1 charge, NH3 is neutral ... right?

OpenStudy (abb0t):

Yes, NH\(_3\) is neutral.

OpenStudy (anonymous):

So if i represent oxidation number of Cobalt as x then : x-1=-2 ?

OpenStudy (anonymous):

that doesn't give 3 ... Please tell me the error of my ways.

OpenStudy (abb0t):

Shouldn't you have = +2?

OpenStudy (anonymous):

how :O ?

OpenStudy (anonymous):

I thought that Cl had a negative charge? That makes it = -2 right?

OpenStudy (abb0t):

Charge of complex ion - total carge of all ligands so, 2+-[5\(\times\)0+(-1)] = 3+

OpenStudy (anonymous):

I still don't understand the representation. you are saying that x+(5*0)-1=2 ?

OpenStudy (anonymous):

Cl2 is the anion right? So how come it got the positive charge?

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

It actually carries a 2 charge since you have two chlorines = -2 hence, to make it neutral, the two Cl\(_2\).

OpenStudy (anonymous):

so you are saying that we take it to be +2 instead of -2 because we need to make it neutral ?

OpenStudy (abb0t):

yes.

OpenStudy (anonymous):

Oh i finally get it!!!!

OpenStudy (anonymous):

[Co(NH3)5 Cl]Cl2 was my original question since Cl has -1 charge the stuff inside the bracket has to be equal to +2 to make their sum 0

OpenStudy (anonymous):

so we equate x-1=+2 my explanation wasn't proper but I got your point. Thanks a tonne ^_^ @abb0t

OpenStudy (abb0t):

yes.

OpenStudy (abb0t):

the counter ion, in this is Cl\(_2\) is there to keep it electrically neutral.

OpenStudy (anonymous):

NO3 is neutral or ... ? whats the charge?

OpenStudy (abb0t):

1-

OpenStudy (anonymous):

What would be the charge for : \[ZnCl _{4}\]

OpenStudy (abb0t):

The charge or oxidation state?

OpenStudy (anonymous):

I was asking for the over all charge. Any way, I learnt quite a long these days. Thanks for your concern :) @abb0t